# The coarse Baum-Connes conjecture with filtered coefficients and product metric spaces

### v1

@@ 3.2. Localization algebras with filtered coefficients
Yu in [20] defined a notion of the localization algebra whose $K$-theory provides a model for $K$-homology. In this subsection, we consider the localization algebra with filtered coefficients.
@@
The localization algebra of $X$ with filtered coefficients in $A$, denoted by $C^{\ast}_{L,f}(X,A)$, is defined to be the norm closure of the $\ast$-algebra consisting of all bounded and uniformly continuous functions $u:[0,\infty)\rightarrow C^{\ast}_{f}(X,A)$ satisfying that
@@
If a $C^{\ast}$-algebra $A$ is equipped with the trivial filtration, then $C^{\ast}_{L,f}(X,A)$ is the localization algebra of $X$ with coefficients in $A$, also denoted by $C^{\ast}_{L}(X,A)$. In particular, when $A=\mathbb{C}$, the $C^{\ast}$-algebra $C^{\ast}_{L,f}(X,\mathbb{C})$ or $C^{\ast}_{L}(X)$ for short is so called the localization algebra of $X$ (cf. [20]).
Let $g:X\rightarrow Y$ be a uniformly continuous coarse map and $\{\delta_{k}\}_{k\in\mathbb{N}}$ be a deceasing sequence of positive numbers with limit zero. Choose a sequence of Borel covers $\{\mathcal{U}_{k}\}_{k\in\mathbb{N}}$ of $Y$ such that
@@
Then $\ell^{2}(U_{k,i}\cap Z_{Y})\otimes H$ is an infinitely dimensional separable Hilbert space for every set $U_{k,i}\in\mathcal{U}_{k}$. Thus we can define an isometry $V_{k,i}:\ell^{2}(g^{-1}(U_{k,i})\cap Z_{X})\otimes H\rightarrow\ell^{2}(U_{k,i}\cap Z_{Y})\otimes H$ with infinitely dimensional cokernel. By the fourth condition as above, there exists $\{U_{k+1,ij}\}_{1\leq j\leq n}\subseteq\mathcal{U}_{k+1}$ such that $\cup^{n}_{j=1}U_{k+1,ij}=U_{k,i}$. Then we can choose $n$ number of isometries $V_{k+1,ij}:\ell^{2}(g^{-1}(U_{k+1,ij})\cap Z_{X})\otimes H\rightarrow\ell^{2}(U_{k+1,ij}\cap Z_{Y})\otimes H$ with infinitely dimensional cokernel. Let $V_{k+1,i}=\bigoplus^{n}_{j=1}V_{k+1,ij}:\ell^{2}(g^{-1}(U_{k,i})\cap Z_{X})\otimes H\rightarrow\ell^{2}(U_{k,i}\cap Z_{Y})\otimes H$ that is an isometry with infinitely dimensional cokernel. Let $W_{k,i}$ be a unitary from the cokernel of $V_{k,i}$ to the cokernel of $V_{k+1,i}$ and let $W^{\prime}_{k,i}=V_{k+1,i}V^{\ast}_{k,i}+0\oplus W_{k,i}$ which is a unitary on $\ell^{2}(U_{k,i}\cap Z_{Y})$. Let $h:S^{1}\rightarrow[0,1)$ be the Borel inverse function of the continuous function $t^{\prime}\mapsto e^{2\pi it^{\prime}}$. Define $\gamma_{k,i}$ to be a path of isometries by $\gamma_{k,i}(t)=e^{2\pi ith(W^{\prime}_{k,i})}V_{k,i}$. Then $\gamma_{k,i}$ is a $2\pi$-Lipschitz path connecting $V_{k,i}$ with $V_{k+1,i}$ and $\operatorname{supp}(\gamma_{k,i}(t))\subseteq g^{-1}(U_{k,i})\times U_{k,i}$. Define
@@
$V_{g}(s)=\gamma_{k}(s-k)\otimes I:\ell^{2}(Z_{X})\otimes H\otimes H_{A}\rightarrow\ell^{2}(Z_{Y})\otimes H\otimes H_{A},\>\>\textrm{if $s\in[k,k+1]$}.$
Then $V_{g}$ is a bounded and uniformly continuous function satisfying that
@@
Let $g:X\rightarrow Y$ be a uniformly continuous coarse map and $\{\delta_{k}\}_{k\in\mathbb{N}}$ be a deceasing sequence of positive numbers with limit zero, then $V_{g}$ as defined above induces a $\ast$-homomorphism
@@
Moreover, for another bounded and uniformly continuous function $V^{\prime}_{g}$ which maps $[0,\infty)$ to the set of all isometries on $\ell^{2}(Z_{X})\otimes H\otimes H_{A}$ and satisfies the condition (3.2), then $Ad^{\prime}_{g}$ defined by $V^{\prime}_{g}$ induces the same homomorphism as $Ad_{g}$ on $K$-theory.
@@
[20, Definition 3.5] Let $X$, $Y$ be two proper metric spaces and $g$, $h$ be two Lipschitz maps from $X$ to $Y$. A continuous homotopy $F(t,x)$ $(t\in[0,1])$ between $g$ and $h$ is said to be strongly Lipschitz if
@@
[20, Definition 3.6] $X$ is said to be strongly Lipschitz homotopy equivalent to $Y$ if there exist two Lipschitz maps $g:X\rightarrow Y$ and $h:Y\rightarrow X$ such that $hg$ and $gh$ are strongly Lipschitz homotopic to $id_{X}$ and $id_{Y}$, respectively.
@@
By the assumption, there exists two Lipschitz maps $g:X\rightarrow Y$ and $h:Y\rightarrow X$ such that $hg$ and $gh$ are strongly Lipschitz homotopic to $id_{X}$ and $id_{Y}$, respectively. It is sufficient to prove that $Ad_{hg}$ induces the identity map on $K_{\ast}(C^{\ast}_{L,f}(X,A))$ and similar to $Ad_{gh}$. Let $F(t,x)$ be a strongly Lipschitz homotopy such that $F(0,\cdot)=hg$ and $F(1,\cdot)=id_{X}$. Then we can choose a sequence of non-negative numbers $\{t_{i,j}\}_{i,j\geq 0}$ in $[0,1]$ such that
@@
Now we prove the lemma for =1$\ast$=1∗ = 1 and by a suspension argument, we can similarly prove for =0$\ast$=0∗ = 0. Let $[u]\in K_{1}(C^{\ast}_{L,f}(X,A))$, let us assume $u\in C^{\ast}_{L,f}(X,A)^{+}$ and let $W=u\oplus I$. Define the following three unitaries:
@@
Then $a_{i}\in M_{2}(C^{\ast}_{L,f}(X,A,H^{\infty}))^{+}$ since that $\operatorname{prop}(a_{i})\leq(C+1)\operatorname{prop}(u(t))+1/(j+1)$ for $t\in[j,j+1]$ and $V_{i}(t)$ acts on the subspace $H_{A}$ as the identity operator, where $C$ is Lipschitz constant of $F$. Moreover, since $d(F(t_{i,j},x),F(t_{i+1,j},x))\leq 1/(j+1)$, thus $[a_{1}]=[a_{2}]$ in $K_{1}(C^{\ast}_{L,f}(X,A))$ (the proof is similar to the last section of the proof of Lemma 3.6). Let $V:H^{\infty}\rightarrow H^{\infty},\>(h_{1},h_{2},\cdots)\mapsto(0,h_{1},\cdots)$, then $I_{\ell^{2}(Z_{X})}\otimes V\otimes I_{H_{A}}$ is an isometric multiplier of $C^{\ast}_{L,f}(X,A,H^{\infty})$. And
@@
$B_{\varepsilon_{t}}(X_{1})\cap B_{\varepsilon_{t}}(X_{2})\subseteq B_{\delta_{t}}(X_{1}\cap X_{2}),$
@@
Let $X=X_{1}\cup X_{2}$ be a uniformly excisive cover. If there exists $\varepsilon_{0}>0$ such that the $\varepsilon_{0}$-neighborhoods of $X_{1},X_{2},X_{1}\cap X_{2}$ are strongly Lipschitz homotopy equivalent to $X_{1},X_{2},X_{1}\cap X_{2}$, respectively, then there is a Mayer-Vietoris six-term exact sequence
@@
$\operatorname{supp}(u(t))\subseteq\{(x,x^{\prime})\in X\times X:d((x,x^{\prime}),X_{i}\times X_{i})\leq c_{t}\},$
@@
$\operatorname{supp}(u(t+t_{0}))\subseteq B_{\varepsilon_{0}}(X_{i})\times B_{\varepsilon_{0}}(X_{i}),$
for all $t\geq 0$. Let $v(t)=u(t+t_{0})$, then $u$ and $v$ are homotopic equivalent in $C^{\ast}_{L,f}(X,A;X_{i})$ since that $u$ is uniformly continuous. Thus the inclusion map from $C^{\ast}_{L,f}(B_{\varepsilon_{0}}(X_{i}),A)$ to $C^{\ast}_{L,f}(X,A;X_{i})$ induces an isomorphism on the $K$-theory. Then $\tau_{\ast}:K_{\ast}(C^{\ast}_{L,f}(X_{i},A))\rightarrow K_{\ast}(C^{\ast}_{L,f}(X,A;X_{i}))$ is an isomorphism by Lemma 3.17.
On the other hand, $C^{\ast}_{L,f}(X,A;X_{i})$ is a ideal of $C^{\ast}_{L,f}(X,A)$ for $i=1,2$. Moreover, we have $C^{\ast}_{L,f}(X,A;X_{1})+C^{\ast}_{L,f}(X,A;X_{2})=C^{\ast}_{L,f}(X,A)$ and $C^{\ast}_{L,f}(X,A;X_{1})\cap C^{\ast}_{L,f}(X,A;X_{2})=C^{\ast}_{L,f}(X,A;X_{1}\cap X_{2})$. Thus we completed the proof by Lemma 2.4. ∎
@@ 3.3. The coarse Baum-Connes conjecture with filtered coefficients
Before building the coarse Baum-Connes conjecture with filtered coefficients, we need introduce a coarsening space for a metric space by Rips complexes.
Let $(N,d_{N})$ be a locally finite metric space, namely, any ball in $N$ just contains finite elements. For any $k>0$, the Rips complex of $N$ at scale $k$, denoted by $P_{k}(N)$, is a simplicial complex whose set of vertices is $N$ and a subset $\{z_{0},\cdots,z_{n}\}\subseteq N$ spans an $n$-simplex in paper $P_{k}(N)$ if and only if $d_{N}(z_{i},z_{j})\leq k$ for any $i,j=0,\cdots,n$.
@@
where $t_{i},t^{\prime}_{i}\in[0,1]$ and $\sum_{i=0}^{n}t_{i}=\sum_{i=0}^{n}t^{\prime}_{i}=1$, and the metric $d_{S_{k}}$ between different connected components is defined to be infinity. Unfortunately, the natural include map from $(N,d_{N})$ to $(P_{k}(N),d_{S_{k}})$ is generally not a coarsely equivalent map. For this reason, we now introduce the following modified metric on $P_{k}(N)$ coming from [18, Definition 7.2.8].
@@
satisfying $ad_{i_{k_{1}k_{2}},\ast}\circ e_{\ast}=e_{\ast}\circ Ad_{i_{k_{1}k_{2}},\ast}$.
@@
By Zorn’s lemma, any proper metric space admits a locally finite net. And any net $N_{X}$ is coarsely equivalent to $X$.
Now, we are ready to show the coarse Baum-Connes conjecture with filtered coefficients in a filtered $C^{\ast}$-algebra $A$ as follow:
@@
Let $X$ be a proper metric space and $N_{X}$ be a locally finite net in $X$, then the following homomorphisms
@@
are isomorphic between abelian groups.
@@
    If $A$ is equipped with the trivial filtration, namely, $A_{r}=A$ for all $r>0$, then the conjecture with filtered coefficients in $A$ is also called the coarse Baum-Connes conjecture with coefficients in $A$. In particular, the coarse Baum-Connes conjecture with filtered coefficients in $\mathbb{C}$ is so called the coarse Baum-Connes conjecture (cf. [8], [19]).
At the last of this section, we consider a uniform version of the coarse Baum-Connes conjecture with filtered coefficients.
@@
Then $C_{f}^{\ast}(\sqcup_{\mathbb{N}}X_{i},A)$ is the norm closure in $\prod_{i\in\mathcal{I}}C_{f}^{\ast}(X_{i},A)$ of the set
$\bigcup_{r>0}\left(\prod_{i\in\mathbb{N}}C_{f}^{\ast}(X_{i},A)_{r}\right).$
And the localization algebra $C_{L,f}^{\ast}(\sqcup_{\mathbb{N}}X_{i},A)$ is the norm closure of the $\ast$-algebra consisting of all bounded and uniformly continuous functions $u:[0,\infty)\rightarrow C_{f}^{\ast}(\sqcup_{\mathbb{N}}X_{i},A)$ satisfying that
$\lim_{t\rightarrow\infty}\sup_{i\in\mathbb{N}}(\operatorname{prop}(u_{i}(t)))=0.$
@@
We call the following conjecture as the uniform coarse Baum-Connes conjecture with filtered coefficients in $A$.
@@
Let $X$ be a proper metric space and $N_{X}$ be a locally finite net in $X$, then the following homomorphisms
@@
are isomorphic between abelian groups, namely, the coarse Baum-Connes conjecture with filtered coefficients in $A$ holds for $\sqcup_{\mathbb{N}}X$.
@@
Actually, the uniform coarse Baum-Connes conjecture with coefficients in $A$ equipped with the trivial filtration is equivalent to the quantitative coarse Baum-Connes conjecture with coefficients in $A$, please refer to [23] for more details.

### v2

@@ 3.2. Localization algebras with filtered coefficients
Yu defined a notion of the localization algebra whose $K$-theory provides a model for $K$-homology in [23]. In this subsection, we introduce the localization algebra with filtered coefficients.
@@
The localization algebra of $X$ with filtered coefficients in $A$, denoted by $C^{\ast}_{L,f}(X,A)$, is defined to be the norm closure of the $\ast$-algebra consisting of all bounded and uniformly continuous functions $u:[0,\infty)\rightarrow C^{\ast}_{f}(X,A)$ satisfying
@@
If a $C^{\ast}$-algebra $A$ is endowed with the trivial filtration, then $C^{\ast}_{L,f}(X,A)$ is the localization algebra of $X$ with coefficients in $A$, also denoted by $C^{\ast}_{L}(X,A)$. In particular, the $C^{\ast}$-algebra $C^{\ast}_{L,f}(X,\mathbb{C})$ or $C^{\ast}_{L}(X)$ for short, is so-called the localization algebra of $X$ (cf. [23]).
Let $g:X\rightarrow Y$ be a uniformly continuous coarse map and $\{\delta_{k}\}_{k\in\mathbb{N}}$ be a decreasing sequence of positive numbers with limit zero. Choosing a sequence of Borel covers $\{\mathcal{U}_{k}\}_{k\in\mathbb{N}}$ of $Y$ such that
@@
Then $\ell^{2}(U_{k,i}\cap Z_{Y})\otimes H$ is an infinite dimensional separable Hilbert space for every set $U_{k,i}\in\mathcal{U}_{k}$. Thus we can define an isometry $V_{k,i}:\ell^{2}(g^{-1}(U_{k,i})\cap Z_{X})\otimes H\rightarrow\ell^{2}(U_{k,i}\cap Z_{Y})\otimes H$ with infinite dimensional cokernel. By the fourth condition as above, there exists $\{U_{k+1,ij}\}_{1\leq j\leq n}\subseteq\mathcal{U}_{k+1}$ such that $\cup^{n}_{j=1}U_{k+1,ij}=U_{k,i}$. Then we can choose a family of isometries $V_{k+1,ij}:\ell^{2}(g^{-1}(U_{k+1,ij})\cap Z_{X})\otimes H\rightarrow\ell^{2}(U_{k+1,ij}\cap Z_{Y})\otimes H$ with infinite dimensional cokernel. Let $V_{k+1,i}=\bigoplus^{n}_{j=1}V_{k+1,ij}:\ell^{2}(g^{-1}(U_{k,i})\cap Z_{X})\otimes H\rightarrow\ell^{2}(U_{k,i}\cap Z_{Y})\otimes H$ which is an isometry with infinite dimensional cokernel. Let $W_{k,i}$ be a unitary from the cokernel of $V_{k,i}$ to the cokernel of $V_{k+1,i}$ and let $W^{\prime}_{k,i}=V_{k+1,i}V^{\ast}_{k,i}+0\oplus W_{k,i}$, which is a unitary on $\ell^{2}(U_{k,i}\cap Z_{Y})$. Let $h:S^{1}\rightarrow[0,1)$ be the Borel inverse function of the continuous function $t^{\prime}\mapsto e^{2\pi it^{\prime}}$. Define $\gamma_{k,i}$ to be a path of isometries by $\gamma_{k,i}(t)=e^{2\pi ith(W^{\prime}_{k,i})}V_{k,i}$. Then $\gamma_{k,i}$ is a $2\pi$-Lipschitz path connecting $V_{k,i}$ with $V_{k+1,i}$ and $\operatorname{supp}(\gamma_{k,i}(t))\subseteq g^{-1}(U_{k,i})\times U_{k,i}$. Define
@@
$V_{g}(s)=\gamma_{k}(s-k)\otimes I_{H_{A}}:\ell^{2}(Z_{X})\otimes H\otimes H_{A}\rightarrow\ell^{2}(Z_{Y})\otimes H\otimes H_{A},\>\>\textrm{if $s\in[k,k+1]$}.$
Then $V_{g}$ is a bounded and uniformly continuous function satisfying
@@
Let $g:X\rightarrow Y$ be a uniformly continuous coarse map and $\{\delta_{k}\}_{k\in\mathbb{N}}$ be a decreasing sequence of positive numbers with limit zero. Then $V_{g}$ as defined above induces a $\ast$-homomorphism
@@
Moreover, for another bounded and uniformly continuous function $V^{\prime}_{g}$ that maps $[0,\infty)$ to the set of all isometries on $\ell^{2}(Z_{X})\otimes H\otimes H_{A}$ and satisfies the condition (3.2), then $Ad^{\prime}_{g}$ defined by $V^{\prime}_{g}$ induces the same homomorphism as $Ad_{g}$ on $K$-theory.
@@
[23, Definition 3.5] Let $X$, $Y$ be two proper metric spaces and $g$, $h$ be two Lipschitz maps from $X$ to $Y$. A continuous homotopy $F(t,x)$ $(t\in[0,1])$ between $g$ and $h$ is said to be strongly Lipschitz if
@@
[23, Definition 3.6] $X$ is said to be strongly Lipschitz homotopy equivalent to $Y$ if there exist two Lipschitz maps $g:X\rightarrow Y$ and $h:Y\rightarrow X$ such that $hg$ and $gh$ are strongly Lipschitz homotopic to $id_{X}$ and $id_{Y}$, respectively.
@@
By the assumption, there exists two Lipschitz maps $g:X\rightarrow Y$ and $h:Y\rightarrow X$ such that $hg$ and $gh$ are strongly Lipschitz homotopic to $id_{X}$ and $id_{Y}$, respectively. It is sufficient to prove that $Ad_{hg}$ induces the identity map on $K_{\ast}(C^{\ast}_{L,f}(X,A))$ and similar to $Ad_{gh}$. Let $F(t,x)$ be a strongly Lipschitz homotopy with the Lipschitz constant $C$ such that $F(0,\cdot)=hg$ and $F(1,\cdot)=id_{X}$. Then we can choose a sequence of non-negative numbers $\{t_{i,j}\}_{i,j\geq 0}$ in $[0,1]$ such that
@@
Now we prove the lemma for =1$\ast$=1∗ = 1 and by a suspension argument, we can similarly prove it for =0$\ast$=0∗ = 0. For an element $[u]\in K_{1}(C^{\ast}_{L,f}(X,A))$, we assume $u\in C^{\ast}_{L,f}(X,A)^{+}$. Let $W=u\oplus I$. Define the following three unitaries:
@@
Since $\operatorname{prop}(a_{i})\leq(C+1)\operatorname{prop}(u(t))+1/(j+1)$ for $t\in[j,j+1]$ and $V_{i}(t)$ acts on $H_{A}$ as the identity operator, thus $a_{i}\in M_{2}(C^{\ast}_{L,f}(X,A,H^{\infty}))^{+}$. Moreover, since $d(F(t_{i,j},x),F(t_{i+1,j},x))\leq 1/(j+1)$, thus $[a_{1}]=[a_{2}]$ in $K_{1}(C^{\ast}_{L,f}(X,A))$ (the proof is similar to the last section of the proof of Lemma 3.6). Let
$V:H^{\infty}\rightarrow H^{\infty},\>(h_{1},h_{2},\cdots)\mapsto(0,h_{1},\cdots),$
then $I_{\ell^{2}(Z_{X})}\otimes V\otimes I_{H_{A}}$ is an isometric multiplier of $C^{\ast}_{L,f}(X,A,H^{\infty})$. And
@@
$B_{\varepsilon_{t}}(X_{1})\cap B_{\varepsilon_{t}}(X_{2})\subseteq B_{\delta_{t}}(X_{1}\cap X_{2})$
@@
Let $X=X_{1}\cup X_{2}$ be a uniformly excisive cover. If there exists $\varepsilon_{0}>0$ such that the $\varepsilon_{0}$-neighborhoods of $X_{1},X_{2}$ and $X_{1}\cap X_{2}$ are strongly Lipschitz homotopy equivalent to $X_{1},X_{2}$ and $X_{1}\cap X_{2}$, respectively. Then there is a Mayer-Vietoris six-term exact sequence
@@
$\operatorname{supp}(u(t))\subseteq\{(x,x^{\prime})\in X\times X:d((x,x^{\prime}),X_{i}\times X_{i})\leq c_{t}\}$
@@
$\operatorname{supp}(u(t+t_{0}))\subseteq B_{\varepsilon_{0}}(X_{i})\times B_{\varepsilon_{0}}(X_{i})$
for all $t\geq 0$. Let $v(t)=u(t+t_{0})$, then $u$ and $v$ are homotopic equivalent in $C^{\ast}_{L,f}(X,A;X_{i})$ since $u$ is uniformly continuous. Thus the inclusion map from $C^{\ast}_{L,f}(B_{\varepsilon_{0}}(X_{i}),A)$ to $C^{\ast}_{L,f}(X,A;X_{i})$ induces an isomorphism on the $K$-theory. Then $\tau_{\ast}:K_{\ast}(C^{\ast}_{L,f}(X_{i},A))\rightarrow K_{\ast}(C^{\ast}_{L,f}(X,A;X_{i}))$ is an isomorphism by Lemma 3.17.
On the other hand, $C^{\ast}_{L,f}(X,A;X_{i})$ is an ideal of $C^{\ast}_{L,f}(X,A)$ for $i=1,2$. Moreover, we have $C^{\ast}_{L,f}(X,A;X_{1})+C^{\ast}_{L,f}(X,A;X_{2})=C^{\ast}_{L,f}(X,A)$ and $C^{\ast}_{L,f}(X,A;X_{1})\cap C^{\ast}_{L,f}(X,A;X_{2})=C^{\ast}_{L,f}(X,A;X_{1}\cap X_{2})$. Thus we completed the proof by Lemma 2.4. ∎
@@ 3.3. The coarse Baum-Connes conjecture with filtered coefficients
Before building the coarse Baum-Connes conjecture with filtered coefficients, we need introduce a class of coarsening spaces for metric spaces by Rips complexes.
Let $(N,d_{N})$ be a locally finite metric space, namely, any ball in $N$ just contains a finite number of elements. For any $k>0$, the Rips complex of $N$ at scale $k$, denoted by $P_{k}(N)$, is a simplicial complex whose set of vertices is $N$ and a subset $\{z_{0},\cdots,z_{n}\}\subseteq N$ spans an $n$-simplex in $P_{k}(N)$ if and only if $d_{N}(z_{i},z_{j})\leq k$ for any $i,j=0,\cdots,n$.
@@
where $t_{i},t^{\prime}_{i}\in[0,1]$ and $\sum_{i=0}^{n}t_{i}=\sum_{i=0}^{n}t^{\prime}_{i}=1$. The metric $d_{S_{k}}$ between different connected components is defined to be infinity. Unfortunately, the natural include map from $(N,d_{N})$ to $(P_{k}(N),d_{S_{k}})$ is generally not a coarsely equivalent map. For this reason, we need introduce the following modified metric on $P_{k}(N)$ coming from [21, Definition 7.2.8].
@@
that satisfies $ad_{i_{k_{1}k_{2}},\ast}\circ e_{\ast}=e_{\ast}\circ Ad_{i_{k_{1}k_{2}},\ast}$.
@@
By Zorn’s lemma, any proper metric space admits a locally finite net. Moreover, any net $N_{X}$ is coarsely equivalent to $X$.
Now, we are ready to introduce the coarse Baum-Connes conjecture with filtered coefficients in a filtered $C^{\ast}$-algebra $A$ as follow:
@@
Let $X$ be a proper metric space and $N_{X}$ be a locally finite net in $X$, then the following homomorphism
@@
is an isomorphism between abelian groups.
@@
    If $A$ is equipped with the trivial filtration, namely, $A_{r}=A$ for each $r>0$. Then the conjecture with filtered coefficients in $A$ is also called the coarse Baum-Connes conjecture with coefficients in $A$. In particular, the coarse Baum-Connes conjecture with filtered coefficients in $\mathbb{C}$ is so-called the coarse Baum-Connes conjecture (cf. [9], [22]).
At the last of this section, we consider a consistent version of the coarse Baum-Connes conjecture with filtered coefficients.
@@
Then the Roe algebra $C_{f}^{\ast}(\sqcup_{\mathbb{N}}X_{i},A)$ is the norm closure in $\prod_{i\in\mathbb{N}}C_{f}^{\ast}(X_{i},A)$ of the set
$\bigcup_{r>0}\left(\prod_{i\in\mathbb{N}}C_{f}^{\ast}(X_{i},A)_{r}\right),$
where $(C_{f}^{\ast}(X_{i},A)_{r})_{r>0}$ is a filtration of Roe algebra $C_{f}^{\ast}(X_{i},A)$ defined in Lemma 3.9. Thus, the localization algebra $C_{L,f}^{\ast}(\sqcup_{\mathbb{N}}X_{i},A)$ is the norm closure of the $\ast$-algebra consisting of all bounded and uniformly continuous functions $u:[0,\infty)\rightarrow C_{f}^{\ast}(\sqcup_{\mathbb{N}}X_{i},A)$ satisfying
$\lim_{t\rightarrow\infty}\sup_{i\in\mathbb{N}}(\operatorname{prop}(u_{i}(t)))=0,$
@@
We call the following conjecture as the consistent coarse Baum-Connes conjecture with filtered coefficients in $A$.
@@
Let $X$ be a proper metric space and $N_{X}$ be a locally finite net in $X$, then the following homomorphism
@@
is an isomorphism between abelian groups, namely, the coarse Baum-Connes conjecture with filtered coefficients in $A$ holds for $\sqcup_{\mathbb{N}}X$.
@@
Actually, the consistent coarse Baum-Connes conjecture with coefficients in $A$ is equivalent to the quantitative coarse Baum-Connes conjecture with coefficients in $A$. Please refer to [26] for more details.